OpenStudy (anonymous):

add: x+4/x^2+x-30+x+2/x^2-x-42

OpenStudy (anonymous):

\[\frac{2 \left(x^3-36 x^2+3\right)}{x^2} \]

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

\[\frac{2 \left(x^3-36 x^2+3\right)}{x^2}=\frac{6}{x^2}+2 x-72 \]